OpenStudy (anonymous):

Could anyone give me assistance on finding all solutions to this particular equation. cos^2x + 2 cos x + 1 = 0 A detailed step by step assistance would be lovely

OpenStudy (zehanz):

Factor it first; if you look closely, you'll see the form a²+2ab+b².

OpenStudy (anonymous):

So it would be (cos + 1)(cos + 2)? It's been a while since I've factored

OpenStudy (anonymous):

go to this website for step by step how to slove this http://www.enotes.com/math/q-and-a/can-solve-equation-like-quadratic-2cos-2x-cos-x-1-212919/

OpenStudy (anonymous):

Lovely must appreciated!

OpenStudy (anonymous):

I would walk you through this problem, but on the internet as you see will very long to solve that is why I found a website with the answer worked out for you.

OpenStudy (anonymous):

toadsage did the site work for you?

OpenStudy (mertsj):

\[(\cos x +1)(\cos x +1)=0\] \[\cos x = -1\] \[x= \pi + 2\pi n\]